8. (10 points) The following information relates to the Windsor Company Ending Inventory (End-of-Year Prices) Index Price Dat8. (10 points) The following information relates to the Windsor Company.DateEnding Inventory(End-of-Year Prices)PriceIndexDecember 31, 2013$ 63,500100December 31, 2014105,434119December 31, 2015113,446131December 31, 2016128,792136December 31, 2017118,158141Use the dollar-value LIFO method to compute the ending inventory for Windsor Company for 2013 through 2017.
Answer #1

Calculate ending inventory

Ending inventory
2013 63500
2014 105434/119 = 88600 (63500+25100*1.19) = 93369
2015 113446/1.31 = 86600 (63500+23100*1.19) = 90989
2016 128792/1.36 = 94700 (90989+8100*1.36) = 102005
2017 118158/1.41 = 83800 63500+20300*1.19 = 87657